Solve the following differential equation:



1. (D²+D)y=sin x



2. (D²+4D+5) y= 50x + 13e^3x



3. (D³+D²-4D-4)y= 4 sin x



4.(D³-D)y=x



5. (D²-4D+4)y=e^x

1
Expert's answer
2021-12-14T10:03:48-0500

1.

"k^2+1=0"

"k=\pm i"

"y_h=c_1cosx+c_2sinx"


"y_p=Axcosx+Bxsinx"

"y'_p=Acosx-Axsinx+Bsinx+Bxcosx=(A+Bx)cosx+(B-Ax)sinx"

"y''_p=Bcosx-(A+Bx)sinx+(B-Ax)cosx-Asinx"


"Bcosx-(A+Bx)sinx+(B-Ax)cosx-Asinx+Axcosx+Bxsinx=sinx"

"B=0"

"-A-A=1"

"A=-1\/2"


"y=y_h+y_p=c_1cosx+c_2sinx-xcosx\/2"


2.

"k^2+4k+5=0"

"k=\frac{-4\pm\sqrt{16-20}}{2}=-2\pm 2i"


"y_h=e^{-2x}(c_1cos2x+c_2sin2x)"


"y_{p1}=Ax+B"

"4A+5Ax+5B=50x"

"A=10,B=-40\/5=-8"

"y_{p1}=10x-8"


"y_{p2}=Ae^{3x}"

"9Ae^{3x}+12e^{3x}+5e^{3x}=13e^{3x}"

"A=-4\/9"

"y_{p2}=-4e^{3x}\/9"


"y=e^{-2x}(c_1cos2x+c_2sin2x)+10x-8-4e^{3x}\/9"


3.

"k^3+k^2-4k-4=0"

"k(k^2-4)+k^2-4=0"

"k_1=-1,k_2=-2,k_3=2"


"y_h=c_1e^{-x}+c_1e^{-2x}+c_1e^{2x}"


"y_p=Acosx+Bsinx"

"y'_p=Bcosx-Asinx"

"y''_p=-Bsinx-Acosx"

"y'''_p=-Bcosx+Asinx"


"-Bcosx+Asinx -Bsinx-Acosx-4(Bcosx-Asinx)-4(Acosx+Bsinx)="

"=4 sin x"

"-5B-5A=0"

"5A-5B=4"

"10A=4\implies A=2/5,B=-2/5"


"y=c_1e^{-x}+c_1e^{-2x}+c_1e^{2x}+2(cosx-sinx)\/5"


4.

"k^3-1=0"

"(k-1)(k^2+k+1)=0"

"k_1=1"

"k^2+k+1=0"

"k_{2,3}=\frac{-1\pm i\sqrt{3}}{2}"


"y_h=c_1e^x+e^{-x/2}(c_2cos(x\sqrt 3/2)+c_3sin(x\sqrt 3/2))"


"y_p=Ax+B"

"-Ax-B=x"

"A=-1,B=0"


"y=c_1e^x+e^{-x/2}(c_2cos(x\sqrt 3/2)+c_3sin(x\sqrt 3/2))-x"


5.

"k^2-4k+4=0"

"k_{1,2}=2"

"y_h=c_1e^{2x}+c_2xe^{2x}"


"y_p=Ae^x"

"A-4A+4A=1"

"A=1"


"y=c_1e^{2x}+c_2xe^{2x}+e^x"